# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 2. Notations and methodology
In the paper [Gol19], the author also claims super approximation for a general group with perfect Zariski closure with respect to $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$. We are reluctant to cite this result unless the group has simple closure, as there seems to be an issue in the paper [SG17] where this claim is based on, and this issue already exists in the case $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$. By first applying sum-product which creates thick one-parameter groups in one simple factor while losing control of the other factor, and then taking commutators, the author essentially arrives at the situation of the aforementioned question for the case $q_{1}=q_{2}=p^{n}$. It appears the author’s argument is that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ is large implies $B$ is large, for which simple counter examples can be found. See the sentence “Lemma 71 and Lemma 34 imply Proposition 37” (Pg. 7254, [SG17]) and trace from there.
To address this question, our approach is as follows. For simplicity we assume $\mathbb{P}_{1}(B)=\Gamma_{q_{1}}$ and $\mathbb{P}_{2}(B)=\Gamma_{q_{2}}$. We consider a section map
@@
such that $\mathbb{P}_{1}\circ\psi$ is the identity map. Take a small parameter $10<\theta<$ 1, e.g. $\theta=\frac{1}{1000}$. According to a dichotomy (Proposition 3.1), there are two scenarios:
@@
    There exists a large exact divisor $q_{3}$ of $q_{2}$, and $x,y\in\Gamma_{q_{1}}$ such that for any $p^{n}||q_{3}$, $\psi(xy)\neq\psi(x)\psi(y)(\operatorname{mod}p^{[n\theta]})$.
@@
    There exists a large exact divisor ${q_{4}}$ of $q_{2}$, and a large subset $S$ of $\Gamma_{q_{1}}$ such that for any $x,y\in S$, any $p^{n}||q_{4}$, $\psi(xy)=\psi(x)\psi(y)(\operatorname{mod}p^{[n\theta]})$.
In the first scenario, one can take $B^{\prime}$ to be the collection of all conjugates of $\psi(xy)\psi(x)^{-1}\psi(y)^{-1}$ by $B$ so that $\mathbb{P}_{1}(B^{\prime})=1$ and $\pi_{q_{3}}\circ\mathbb{P}_{2}(B^{\prime})$ is large, thus $\pi_{q_{1},q_{3}}(BB^{\prime})$ is large in $\Gamma_{q_{1}}\times\Gamma_{q_{3}}$.
In the second scenario, we consider two cases. If for a large exact divisor $q_{4}^{\prime}$ of ${q_{4}}$, we have $\pi_{q_{3}^{\{\theta/2\}}}\circ\mathbb{P}_{2}\circ\psi=1$, then by taking commutator of $\psi(\Gamma_{q_{1}})$ iteratively one can create a set $B^{\prime\prime}$ such that $\pi_{q_{4}^{\prime}}\circ\mathbb{P}_{2}(B^{\prime\prime})=1$ and $\mathbb{P}_{1}(B^{\prime\prime})$ is large in $\Gamma_{q_{1}}$, so that $\pi_{q_{1},q_{4}^{\prime}}(BB^{\prime\prime})$ is large in $\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime}}$.
If for a large exact divisor $q_{4}^{\prime\prime}$ of ${q_{4}}$ such that for any $p^{n}||q_{4}^{\prime\prime}$, we have $\pi_{q_{3}^{\{{\theta}/{2}\}}}\circ\mathbb{P}_{2}\circ\varphi\neq 1$, then we must have $q_{4}^{\prime\prime}||q_{1}$, and we can find a set $B^{\prime\prime\prime}\subset\text{Im}(\psi)$ such that $B^{\prime\prime\prime}$ contains all the preimage of a set $\mathcal{P}\subset\Gamma_{Q_{1}^{2}}\times\Gamma_{Q_{2}^{2}}$ under the following canonical projection:
$\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime\prime}}\cong\Gamma_{q_{1}/q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\rightarrow\Gamma_{q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\rightarrow\Gamma_{Q_{1}^{2}}\times\Gamma_{Q_{2}^{2}}$
where $Q_{1},Q_{2}|q_{4}^{\prime\prime}$ such that for each $p^{n}||q_{4}^{\prime\prime}$, $v_{p}(Q_{1}),v_{p}(Q_{2})=\Theta(n\theta)$, and $\mathcal{P}=\{(1,1)+n(Q_{1}X_{1},Q_{2}X_{2})\}_{n\in\mathbb{Z}}$, where $X_{1},X_{2}$ are two traceless primitive matrices in $\text{Mat}(2,\mathbb{Z})$. Then one can apply quantitative non-concentration of random walks to find elements from $A$ to conjugates $(X_{1},X_{2})$ to different directions. Multiplying these conjugates together and taking commutators, we can recover a large set in $\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime\prime}}$.
This whole scheme is carried out in the proof of Proposition 6.1. One can then apply Proposition 6.1 iteratively to create a large subset of $\Lambda_{q}$, which forces (2.4) to hold afterall.
The idea of considering section maps comes from Bourgain’s gluing scheme for $\mathbb{Z}_{q}$ (See Lemma 6.71 of [Bou08]). If the group is non-semismple, this method can also be applied to connect the semisimple part of the group to the unipotent part. See Section 8.
@@
Since Proposition 5.1 is directly given by the main theorem of [BV12], if only for a proof for Theorem 1.2, one can skip Section 5 and start from Section 6, and reference back to more or less standard lemmas in Sections 3 and 4 when necessary. We include a detailed proof of Proposition 5.1 to deal with groups with denominators (Theorem 1.3) and for potential further generalization.
@@ 3. Preliminaries on combinatorics
Let $G_{1},G_{2}$ be two finite groups and let $\psi:G_{1}\rightarrow G_{2}$ some map. Then for $0<\varepsilon<\frac{1}{1600}$ we have either
@@
Remark 3.4.
If $\text{gcd}\left(\left|G_{1}\right|,\left|G_{2}\right|\right)=1$, then (3.3) implies
|$\left|\left\{x\in G_{1}\mid\psi(x)\neq 1\right\}\right|<\sqrt{\varepsilon}\left|G_{1}\right|.$
@@
Let $A$ be a finite subset of a group $Z$ and $\mathcal{G}\subset A\times A$, $40<\varepsilon<1/$ 4, such that
@@
Proof of Lemma.
3.5. The proof was stated in [Bou08] for an abelian group $Z$. The same proof works for a general group as well. ∎
@@
$A=\left\{(x,\psi(x))\mid x\in G_{1}\right\}\subset G_{1}\times G_{2}$
and let $\mathcal{G}^{\prime}=\{((x,\psi(x)),(y,\psi(y))):(x,y)\in\mathcal{G}\}$. Then $\mathcal{G}^{\prime}$ is a graph on $A$ and $|\mathcal{G}^{\prime}|=|\mathcal{G}|\geq(1-\epsilon)|G_{1}|^{2}=(1-\epsilon)|A|^{2}$. Apply Lemma 3.5 with $Z=G_{1}\times G_{2}$. We obtain a subset $A^{\prime}\subset A$ satisfying
$\left|A^{\prime}\right|>(1-\sqrt{\varepsilon})\left|G_{1}\right|$ (3.8)
@@
Let $H_{1}=\mathbb{P}_{1}\left(H\right)$. We have
[…]
implying that $H_{1}=G_{1}$. Then for any $x\in G_{1},$ there exists $f(x)\in G_{2}$ such that $(x,f(x))\in H$. Assume there exists $(y,z_{1}),(y,z_{2})\in H$ with $z_{1}\neq z_{2}$. Then for any $x\in G_{1}$, we have $(x,f(x)z_{1}z_{2}^{-1})\in H$ with $(x,f(x)z_{1}z_{2}^{-1})\neq(x,f(x))$. So $|H|\geq 2|G_{1}|\geq 2|A^{\prime}|$. Contradiction. Hence the choice of $f(x)$ is unique for all $x$. Since $H$ is a subgroup, we get $(1,1)\in H$ so $f(1)=1$. Also we see $(y_{1},f(y_{1}))(y_{2}^{-1},f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1})f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1}y_{2}^{-1}))$, so $f$ is a group homomorphism.
@@
by inclusion-exclusion; and hence
@@
Therefore, $\mathbb{P}_{1}\left(A^{\prime}\right)\cap\left[\mathbb{P}_{1}\left(A^{\prime}\right)\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}\mathbb{P}_{1}\left(A^{\prime}\right)\right]\neq\emptyset$, and for $\left(y_{1},y_{2}\right)\in\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)$ with $y_{1}y_{2}\in\mathbb{P}_{1}\left(A^{\prime}\right)$, we get
$x_{2}f(x_{1}^{-1}y_{1}y_{2})=\psi\left(y_{1}y_{2}\right)=\psi\left(y_{1}\right)\psi\left(y_{2}\right)=x_{2}f(x_{1}^{-1}y_{1})x_{2}f(x_{1}^{-1}y_{2}).$
@@
We also need the following bounded generation result over $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$.
@@
In [Hel08] Helfgott proved that given $A\subset\text{SL}_{2}(\mathbb{F}_{p})$ with $|A|>p^{\frac{8}{3}}$, then $A\cdot A\cdot A={\text{SL}}_{2}(\mathbb{F}_{p})$ for a sufficiently large prime $p$. Helfgott’s result can be obtained in an elegant way by a representation theoretical approach due to Gowers [Gow08], but Gower’s approach does not seem to adapt well for a general composite modulus. We follow Helfgott’s approach for the proof of Proposition 3.12.
@@
. Since $|A|>(q_{1}q_{2})^{3-\delta}$, by the pigeon hole principle, there exists $\vec{v}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{2},\vec{w}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$, such that the cardinality of the set
$\{(\gamma_{1},\gamma_{2}):\gamma_{1}\text{ has lower row }\vec{v},\gamma_{2}\text{ has lower row }\vec{w}\}$
@@
Define an equivalence relation $\sim$ on $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q_{2}\mathbb{Z})$ as:
$(\gamma_{1},\gamma_{2})\sim(\gamma_{1}^{\prime},\gamma_{2}^{\prime})$
if and only if the second rows of $\gamma_{1}$, $\gamma_{1}^{\prime}$ are the same up to a scaler in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$, and the second rows of $\gamma_{2}$ $\gamma_{2}^{\prime}$ are the same up to a scaler in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$. There are at most $q_{1}q_{2}$ many such classes. By pigeon hole, there exists one class which contains at least $(q_{1}q_{2})^{2-\delta}$ many elements from $A$. This implies the set
@@
has cardinality $>(q_{1}q_{2})^{2-\delta}$. By pigeon hole again, there exists $x_{0}\in\mathbb{Z}/q_{1}\mathbb{Z},y_{0}\in\mathbb{Z}/q_{2}\mathbb{Z}$ such that
@@
$\left(\begin{matrix}\lambda&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)\left(\begin{matrix}1&x\\ 0&1\end{matrix}\right)\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)^{-1}=\left(\begin{matrix}1&\lambda^{2}x\\ 0&1\end{matrix}\right).$ (3.18)
Using (3.18) and applying Lemma 3.14 to the set

### v4

@@ 2. Notations and methodology
We need to consider the case $q_{1},q_{2}$ coprime for Question 2.7 when growing $q_{*}$ to a very large divisor of $q$ for one simple factor. We also need to consider $q_{1},q_{2}$ non-coprime when passing from one simple factor to another. In the paper [SG17], the author claims super approximation for a general group with semisimple closure with respect to $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$. There is an issue, which already exists in the case $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$. In the proof, The author arrives at the situation of Question 2.7 for $q_{1}=q_{2}=p^{n}$. It appears the author’s argument is that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ being large implies $B$ is large, for which simple counter examples can be found (e.g. diagonal groups). This issue can be resolved if we also allow some extra help from the set $A$ given by Proposition 2.2. This is good enough for proving Proposition 2.2.
We develop a gluing tool (Proposition 6.1) to address a more general situation of Question 2.7, which in our view is a main technical contribution of this paper. We give a (more detailed) sketch of solution to Question 2.7 here, which contains all essential ideas for Proposition 6.1. Related to our application, we assume $n_{i}$ is large for each $i\in I_{2}$ (say, $n_{i}>\theta^{-2}$ for a parameter $\theta$ introduced below), and assume $\mathbb{P}_{2}(B)\subset\Gamma(\tilde{q}_{2})/\Gamma(q_{2})$, where $\tilde{q}_{2}$ is the square free part of $q_{2}$, so that $\Gamma(\tilde{q}_{2})/\Gamma(q_{2})$ is a product of $p$-groups, and the loss of size from $\Gamma_{q_{2}}$ to $\Gamma(\tilde{q}_{2})/\Gamma(q_{2})$ is insignificant. We do not have any such requirement for $q_{1}$, e.g. $q_{1}$ can be square free.
Without hiding key features, for simplicity we assume $\mathbb{P}_{1}(B)=\Gamma_{q_{1}}$ and $\mathbb{P}_{2}(B)=\Gamma(\tilde{q}_{2})/\Gamma({q_{2}})$, where $\tilde{q}_{2}$ is the square free part of $q_{2}$. Our goal is to show that a bounded product of $B\cup A$ covers a large subset of $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$.
Consider a section map
@@
such that $\mathbb{P}_{1}\circ\psi$ is the identity map. Take a small parameter $0<\theta<1$, e.g. $\theta=10^{-12}$. For each $i\in I_{2}$ (i.e. $p_{i}^{n_{i}}\|q_{2}$), we consider the map $\psi_{i}=\pi_{p^{[n\theta]}}\circ\mathbb{P}_{2}\circ\psi$. According to a dichotomy (Proposition 3.1), there are two scenarios:
@@
    There exists $\mathcal{G}_{j}\subset\Gamma_{q_{1}}\times\Gamma_{q_{1}},|\mathcal{G}_{j}|>10^{-4}|\Gamma_{q_{1}}|^{2}$, such that $\psi_{j}(xy)\neq\psi_{j}(x)\psi_{j}(y)$, $\forall(x,y)\in\mathcal{G}_{j}$.
@@
    There exists $S_{j}\subset\Gamma_{q_{1}},|S_{j}|\geq\frac{99}{100}|\Gamma_{q_{1}}|$, such that over $S_{j}$, $\psi_{j}$ agrees with a homomorphism $h_{j}:\Gamma_{q_{1}}\rightarrow\Gamma_{p^{[n\theta]}}$.
Split $I_{2}=J_{1}\sqcup J_{2}$, where $J_{1}$ is the collection of indices when Case (1) happens, and $J_{2}$ is the complement. We further split $J_{2}=J_{3}\sqcup J_{4}$ where $J_{3}$ is the set of indices such that $h_{j}$ is trivial at the half level, i.e. $\pi_{p_{j}^{[\frac{1}{2}{n_{j}\theta}]}}\circ h_{j}=1$, and $J_{4}$ is the complement.
Case 1: $q^{J_{1}}\geq(q_{2})^{\frac{1}{2}}$. A probability argument shows that there is a large subset $\mathcal{G}\subset\Gamma_{q_{1}}\times\Gamma_{q_{1}}$ (in fact $|\mathcal{G}|>q_{2}^{0-}|\Gamma_{q_{1}}|^{2}$), and $U_{0}\subset J_{1}$ with $q^{U_{0}}\geq(q^{J_{1}})^{\frac{1}{2}10^{-4}}$, such that $\psi_{j}(g_{1})\psi_{j}(g_{2})\neq\psi_{j}(g_{1}g_{2}),\forall(g_{1},g_{2})\in\mathcal{G},\forall j\in U_{0}$. Take any $(g_{1},g_{2})\in\mathcal{G}$. Let $w=\psi(g_{1})\psi(g_{2})(\psi(g_{1}g_{2}))^{-1}$, and $w_{B}=\{[[w,b_{1}],b_{2}]:b_{1},b_{2}\in B\}$. From Lemma 3.39 and Lemma 3.33, we have $\mathbb{P}_{1}((w_{B})^{N_{1}})=1$ and $\pi_{q^{U_{0}}}\circ\mathbb{P}_{2}((w_{B})^{N_{1}})$ contains a large congruence subgroup of $\Gamma_{q^{U_{0}}}$, for some $N_{1}=O(\log\frac{1}{\theta})$. Therefore, $\pi_{q_{1},q^{U_{0}}}(B\cdot(w_{B})^{N_{1}})$ covers a large subset of $\Gamma_{q_{1}}\times\Gamma_{q^{U_{0}}}$.
Case 2.1: $q^{J_{2}}\geq(q_{2})^{\frac{1}{4}}$. Similar to Case 1, we can find a large subset $W_{1}\subset\Gamma_{q_{1}}$, and $U_{1}\subset J_{2}$ with $q^{U_{1}}\geq(q^{J_{2}})^{\frac{99}{200}}$, such that $\pi_{p_{j}^{[\frac{1}{2}{n_{j}\theta}]}}(g)\circ\psi_{j}=1,\forall g\in W_{1},\forall j\in U_{1}$. Then applying Lemma 3.33, for some $N_{2}=O(\log\frac{1}{\theta})$, $\mathbb{P}_{1}([W_{1},W_{1}]^{N_{2}})$ covers a large congruence subgroup of $\Gamma_{q_{1}}$, while $\pi_{p_{j}^{2[\frac{1}{2}n_{j}\theta]}}\circ\mathbb{P}_{2}([W_{1},W_{1}]^{N_{2}})=1,\forall j\in U_{1}$, noticing that we have increased the exponent of $p_{j}$ by a factor of 2 after taking commutator. Keep iterating until we obtain a set $B_{1}\subset[W_{1},W_{1}]^{O((\log\frac{1}{\theta})^{2})}$ such that $\mathbb{P}_{1}(B_{2})$ remains large in $\Gamma_{q_{1}}$, but $\mathbb{P}_{2}(B_{1})=1(\operatorname{mod}q^{U_{1}})$. Then, $\pi_{q_{1},q^{U_{1}}}(B\cdot B_{1})$ covers a very large subset of $\Gamma_{q_{1}}\times\Gamma_{q^{U_{1}}}$.
Case 2.2: $q^{J_{3}}\geq(q_{2})^{\frac{1}{4}}$. In this case, we can find a large set $W_{2}\subset\Gamma_{q_{1}}$ and $U_{2}\subset J_{3}$, such that over $W_{2}$, $\pi_{p_{j}^{[n_{j}\theta]}}\circ\mathbb{P}_{2}\circ\psi=h_{j},\forall j\in U_{2}$. The set $W_{2}$ boundedly generates a large subgroup $G$ of $\Gamma_{q_{1}}$. The homomorphisms $\{h_{j}:j\in U_{2}\}$ uniquely determine a homomorphism $h:\Gamma_{q_{1}}\rightarrow\Gamma_{{(q^{U_{2}})^{\{\theta\}}}}$ such that $\pi_{p_{j}^{[n_{j}\theta]}}\circ h=h_{j},\forall j\in U_{2}$. Thus we can construct a section map
$\bar{\psi}:G\rightarrow W_{2}^{O(1)},$
such that $\mathbb{P}_{1}\circ\bar{\psi}$ is the identity map, and $\pi_{(q^{U_{2}})^{\{\theta\}}}\circ\mathbb{P}_{2}\circ\bar{\psi}=h$. The homomorphism $h$ is completely determined by homomorphisms between local factors of $G$ and $\Gamma_{(q^{U_{2}})^{\{\theta\}}}$. Also recall $\mathbb{P}_{2}(B)=\Gamma({\tilde{q}_{2}})/\Gamma(q_{2})$, which implies $h_{j}(G)$ is a $p_{j}$-group for each $j\in U_{2}$. Thus according to Lemma 3.44, since $h_{j}$ is nontrivial for each $j\in U_{2}$, we must have $j\in I_{1}$, so $U_{2}\subset I_{1}$. By working locally first and then applying Chinese Remainder Theorem, we can find an element $g_{0}\in G$ such that $\pi_{q^{I_{1}-U_{2}}}\circ\mathbb{P}_{1}(g_{0})=1$ and $\pi_{p_{j}^{[\frac{1}{2}n_{j}\theta]}}\circ\psi_{j}(g_{0})\neq 1,\forall j\in U_{2}$. Lemma 3.44 implies $v_{p_{j}}(g_{0}-1)=O(\theta n_{j}),\forall j\in U_{2}$ (More precisely, this estimate holds for almost all $j\in U_{2}$, with exceptions coming from those $j$ where $\pi_{p_{j}^{n_{j}}}(G)$ is not large, which are not many since $G$ is a large subgroup of $\Gamma_{q_{1}}$. Let us ignore this minor issue here).
Let $g=g_{0}^{(q^{U_{2}})^{\{\frac{1}{4}\theta\}}}$. Then $\{\bar{\psi}(g^{n}):n\in\mathbb{N}\}$ satisfies
$\pi_{q^{I_{1}-U_{2}}}\circ\mathbb{P}_{1}(\bar{\psi}(g^{n}))=1$
$\bar{\psi}(g^{n})\equiv(1+nQ_{1}X_{1},nQ_{2}X_{2})(\operatorname{mod}Q_{1}^{2},Q_{3})$ (2.9)
where $X_{1},X_{2}\in\text{Mat}_{2}(\mathbb{Z})$ traceless and primitive, $Q_{1},Q_{2}|q^{U_{2}}$ satisfies $v_{p}(Q_{1})=v_{p}(g-1)$, $v_{p}(Q_{2})=v_{p}(\mathbb{P}_{2}\circ\bar{\psi}(g)-1)$, which uniquely determines $Q_{1}$ and $Q_{2}$, and $Q_{3}=\text{gcd}(Q_{2}^{2},(q^{U_{2}})^{\{\theta\}})$, where we need $\pi_{Q_{3}}\circ\mathbb{P}_{2}\circ\bar{\psi}$ is a homomorphism. Importantly, we have the following valuation estimates:
$v_{p_{j}}(Q_{1}),v_{p_{j}}(Q_{2}),v_{p_{j}}(Q_{3}),v_{p_{j}}(Q_{3})-v_{p_{j}}(Q_{2})=\Theta(n_{j}\theta),\hskip 14.22636pt\forall j\in U_{2},$
where the implied constants for $\Theta(\cdot)$ are absolute. This explains our choice of $g$. Then (2.1) gives a “thick” one parameter group whose projection to each component is thick. We can then use random walks (Proposition 4.2) to find $g_{1},g_{2},g_{3},g_{4},g_{5}\in A$ to conjugate $X=(X_{1},X_{2})$ to different directions so that
$\text{gcd}(\text{Det}(X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}),q^{U_{2}})$
is small. Multiplying $\{\bar{\psi}(g^{n}):n\in\mathbb{N}\}$ with all of its $g_{i}$-conjugates together, $1\leq i\leq 5$, and taking a further product set yield a set $B_{2}\subset\{B\cup A\}^{O(\log\frac{1}{\theta})}$ such that $\pi_{q^{I_{1}-U_{2}}}\circ\mathbb{P}_{1}(B_{2})=1$ and $\pi_{q^{U_{2}},q^{U_{2}}}(B_{2})$ is large in $\Gamma_{q^{U_{2}}}\times\Gamma_{q^{U_{2}}}$. Therefore, $\pi_{q_{1},q^{U_{2}}}(B\cdot B_{2})$ is a very large subset in $\Gamma_{q_{1}}\times\Gamma_{q^{U_{2}}}$.
In each of the above case, we manage to find a set from a bounded (depending on $\theta$) product of $B\cup A$ whose reduction by $(q_{1},q^{U_{i}}),0\leq i\leq 2$ is large, compared to the set $B$ whose reduction is known to be large at $(q_{1},1)$. It is noted that the progress of modulus increase, measured by $\frac{\log q^{U_{i}}}{\log q_{2}}$ is bounded below by an absolute positive constant. One can then iterate the above gluing process boundedly many times to improve the modulus of the second component to a very large exact divisor of $q_{2}$, thus giving an affirmative answer to Question 2.7 (with extra help from AA!).
Regarding the proof of Proposition 2.2, write $q=q_{s}q_{l}$ where all exponents of prime factors of $q_{s}$ are small and all exponents of prime factors of $q_{l}$ are large. If $q_{s}$ is not too small, we can first apply Theorem A to show $\pi_{q_{s}}(A)$ is big in $\Gamma_{q_{s}}\times\Gamma_{q_{s}}$, and iteratively apply Proposition 5.1 (which helps to create large sets under smaller moduli) and Proposition 6.1 (which helps to glu) to improve the moduli, until we find a bounded product of $A$ covers a large congruence subgroup of $\Lambda_{q}$ of size at least $|\Lambda_{q}|^{1-\frac{\varepsilon}{2}}$. This forces (2.4) to hold for a sufficiently small choice of $\delta$. If $q_{s}$ is too small, we can just work with $q_{l}$ and directly start the iterative process.
@@
The discussions of Case 1, Case 2.1, Case 2.2 correspond to the analysis in §6.1, §6.2 and §6.3 for the proof of Proposition 6.1. Caution: Some notations appearing above, such as $\mathbb{P}_{1},\mathbb{P}_{2},Q_{1},Q_{2},Q_{3},W_{1},W_{2},G$ have different meanings in §6.
Remark 2.11.
One can start reading from Section 5 and reference back to more or less standard lemmas in Sections 3 and 4 when necessary. Since Proposition 5.1 is directly given by the main theorem of [BV12], if only for a proof for Theorem 1.2, one can assume Proposition 5.1 and skip the rest part in Section 5. We include a detailed proof of Proposition 5.1 to deal with groups with denominators (Theorem 1.3).
@@ 3. Preliminaries on combinatorics
Let $G_{1},G_{2}$ be two finite groups and let $\psi:G_{1}\rightarrow G_{2}$ be some map. Then for $0<\varepsilon<\frac{1}{1600}$ we have either
@@
Let $A$ be a finite subset of a group $Z$ and $\mathcal{G}\subset A\times A$, $0<\varepsilon<1/4$, such that
@@
Proof of Lemma 3.4.
The proof was stated in [Bou08] for an abelian group $Z$. The same proof works for a general group as well. ∎
@@
$A=\left\{(x,\psi(x))\mid x\in G_{1}\right\}\subset G_{1}\times G_{2},$
and
$\mathcal{G}^{\prime}=\{((x,\psi(x)),(y,\psi(y))):(x,y)\in\mathcal{G}\}.$
Then $|\mathcal{G}^{\prime}|=|\mathcal{G}|\geq(1-\varepsilon)|G_{1}|^{2}=(1-\varepsilon)|A|^{2}$. Applying Lemma 3.4 with $Z=G_{1}\times G_{2}$ to the set $A$, we obtain a subset $A^{\prime}\subset A$ satisfying
$\left|A^{\prime}\right|>(1-\sqrt{\varepsilon})\left|A\right|=(1-\sqrt{\varepsilon})\left|G_{1}\right|$ (3.7)
@@
Let $\mathbb{P}_{1}$ be the projection map $G_{1}\times G_{2}\rightarrow G_{1}$. Let $H_{1}=\mathbb{P}_{1}\left(H\right)$. We have
@@
implying that $H_{1}=G_{1}$. Then for any $x\in G_{1},$ there exists $f(x)\in G_{2}$ such that $(x,f(x))\in H$. Assume there exists $(y,z_{1}),(y,z_{2})\in H$ with $z_{1}\neq z_{2}$. Then for any $x\in G_{1}$, we have $(x,f(x)z_{1}z_{2}^{-1})\in H$ with $(x,f(x)z_{1}z_{2}^{-1})\neq(x,f(x))$. So $|H|\geq 2|G_{1}|\geq 2|A^{\prime}|$. Contradiction. Hence the choice of $f(x)$ is unique for all $x$. Since $H$ is a subgroup, we get $(1,1)\in H$ so $f(1)=1$. Also we see $(y_{1},f(y_{1}))(y_{2}^{-1},f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1})f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1}y_{2}^{-1}))$ by the uniqueness of choice for $f(y_{1}y_{2}^{-1})$, so $f$ is a group homomorphism. Therefore, over $A^{\prime}$ we have
$(x,\psi(x))=(x_{1},x_{2})\cdot(x_{1}^{-1}x,f(x_{1}^{-1}x))\Rightarrow\psi(x)=x_{2}f(x_{1}^{-1}x).$ (3.11)
@@
by inclusion-exclusion. Hence
@@
Therefore, $\mathbb{P}_{1}\left(A^{\prime}\right)\cap\left[\mathbb{P}_{1}\left(A^{\prime}\right)\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}\mathbb{P}_{1}\left(A^{\prime}\right)\right]\neq\emptyset$. For $\left(y_{1},y_{2}\right)\in\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)$ with $y_{1}y_{2}\in\mathbb{P}_{1}\left(A^{\prime}\right)$, we get
[…]
@@
We also need the following bounded generation result over congruence quotients of $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$.
@@
In [Hel08] Helfgott proved that given $A\subset\text{SL}_{2}(\mathbb{F}_{p})$ with $|A|>p^{\frac{8}{3}}$, then $A\cdot A\cdot A={\text{SL}}_{2}(\mathbb{F}_{p})$ for a sufficiently large prime $p$. Helfgott’s result can be obtained in an elegant way by a representation theoretical approach due to Gowers [Gow08], but Gower’s approach does not seem to have an easy generalization for a general composite modulus. We follow Helfgott’s approach for the proof of Proposition 3.12.
@@
Since $|A|>(q_{1}q_{2})^{3-\delta}$, by the pigeon hole principle, there exists $\vec{v}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{2},\vec{w}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$, such that the cardinality of the set
[…]
@@
Next, we define an equivalence relation $\sim$ on $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q_{2}\mathbb{Z})$ as: $(\gamma_{1},\gamma_{2})\sim(\gamma_{1}^{\prime},\gamma_{2}^{\prime})$ if and only if the second rows of $\gamma_{1}$, $\gamma_{1}^{\prime}$ are the same up to a scaler multiple in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$, and the second rows of $\gamma_{2}$ $\gamma_{2}^{\prime}$ are the same up to a scaler multiple in $(\mathbb{Z}/q_{2}\mathbb{Z})^{*}$. There are at most $(q_{1}q_{2})^{1+}$ many such classes. By pigeon hole, there exists one class which contains at least $(q_{1}q_{2})^{2-\delta-}$ many elements from $A$. This implies the set
@@
has cardinality exceeding $(q_{1}q_{2})^{2-\delta-}$. By pigeon hole again, there exists $x_{0}\in\mathbb{Z}/q_{1}\mathbb{Z},y_{0}\in\mathbb{Z}/q_{2}\mathbb{Z}$ such that
@@
$\left(\begin{matrix}\lambda&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)\left(\begin{matrix}1&x\\ 0&1\end{matrix}\right)\left(\begin{matrix}\lambda&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)^{-1}=\left(\begin{matrix}1&\lambda^{2}x\\ 0&1\end{matrix}\right).$ (3.18)
Using (3.18) and applying Lemma 3.14 to the sets